A five-year project has an initial fixed asset investment of $315,000, an initial NWC investment of $25,000, and an annual OCF of −$35,000. The fixed asset is fully depreciated over the life of the project and has no salvage value. If the required return is 11 percent, what is this project’s equivalent annual cost, or EAC? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
1
Expert's answer
2020-10-21T10:17:39-0400

The NPV is:

"NPV = -(315,000 + 25,000) + \frac{-35,000 + 63,000}{1 + 0.11} + \frac{-35,000 + 63,000}{1.11^2} + \frac{-35,000 + 63,000}{1.11^3} + \frac{-35,000 + 63,000}{1.11^4} + \frac{-35,000 + 63,000 + 25,000}{1.11^5} = -221,678.6."

The project’s equivalent annual cost is:

"EAC = \frac{-221,678.6*0.11}{1 - 1.11^{-5}} = -59,979.65."